OpenStudy (anonymous):

BEST RESPONSE Express the complex number in trigonometric form. -3i

OpenStudy (aravindg):

-3(cos 90+i sin 90)

OpenStudy (anonymous):

are you sure?

OpenStudy (aravindg):

yep :) where 90 is in degrees

OpenStudy (anonymous):

I don't have a negative 3 answer choice though, so would it be the same only with a positive 3?

OpenStudy (aravindg):

is this a test ?

OpenStudy (anonymous):

no its just a homework assignment

OpenStudy (aravindg):

ok then u can wrote it also as 3(cos 90- i sin 90)

OpenStudy (anonymous):

Thank you! Can you help me with one more similar one?

terenzreignz (terenzreignz):

No, that won't do... you can't have cos A - isin A

OpenStudy (aravindg):

3(cos 90- i sin 90)=-3 sin 90

terenzreignz (terenzreignz):

Unless it's in the choices, of course, I'd concede :)

OpenStudy (aravindg):

*-3i sin 90

terenzreignz (terenzreignz):

It has to be cos A + i sin A never a difference. It's more like 3(cos 270 + i sin 270)

OpenStudy (anonymous):

so which one is it?

terenzreignz (terenzreignz):

@AravindG What do you think? :>

OpenStudy (aravindg):

ya u r right :)

terenzreignz (terenzreignz):

Yay, I convinced @AravindG Better add that to my list of achievements LOL anyway, yeah, that's it :)

terenzreignz (terenzreignz):

You had another question @danny369 ?

OpenStudy (anonymous):

Yes, Express the complex number in trigonometric form. -3 + 3 squareoot (3) i

terenzreignz (terenzreignz):

I knew this had to happen sometime :D First, acquaint yourself with the complex plane :D For a complex number a + bi 

terenzreignz (terenzreignz):



terenzreignz (terenzreignz):

So which is it?

OpenStudy (anonymous):

the box that is 90- 180 degrees

terenzreignz (terenzreignz):

right :) So, to get the polar form r(cos A + i sin A) You need r, and you need A :D getting r is relatively straightforward, use this formula \[\large r = \sqrt{a^2 + b^2}\]

OpenStudy (aravindg):

just to add in here : \[-\pi < \theta \le \pi \] is often used to repesent \(\theta \) as principal argument

OpenStudy (anonymous):

okay so r would be 4.4426?

OpenStudy (anonymous):

i mean 4.2426

terenzreignz (terenzreignz):

Try it again \[\large a = 3 \ \ ; \ \ b=3\sqrt3\]

OpenStudy (anonymous):

ohh so its 6

terenzreignz (terenzreignz):

Yep. Now finding the angle A is a bit tricky, first you might want to evaluate this \[\large A = \tan^{-1}\frac{b}{a}\]

OpenStudy (anonymous):

is it 60?

terenzreignz (terenzreignz):

Well, almost. But remember, you picked the quadrant where the angle is from 90 - 180. So what's it gonna be? :)

terenzreignz (terenzreignz):

Okay, in general, this is going to be a mouthful, but it's accurate, trust me :) given the complex number a + bi Use these rules If a and b are both positive, then\[\large A = \tan^{-1}\left|\frac{b}{a}\right|\] If a is negative and b is positive\[\large A = 180^o -\tan^{-1}\left|\frac{b}{a}\right|\] If both a and b are negative\[\large A = 180^o +\tan^{-1}\left|\frac{b}{a}\right|\] If a is positive and b is negative \[\large A = 360^o -\tan^{-1}\left|\frac{b}{a}\right|\]

OpenStudy (anonymous):

Im confused

terenzreignz (terenzreignz):

Well, as you pointed out, a is negative and b is positive, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, you got 60 for the angle right? But a is negative and b is positive, if you consult that mouthful of rules I gave you, you'll find that the angle A is equal to 180 - 60 = 120 Get it now? :)

OpenStudy (anonymous):

oh okay, yes I understand what do we do next?

terenzreignz (terenzreignz):

that's it :) r = 6 A = 120 So just plug in \[\large r\left(\cos A + i\sin A\right)\]

terenzreignz (terenzreignz):

Got it yet?

OpenStudy (anonymous):

I understand, but the answers come in the form of pi

terenzreignz (terenzreignz):

oh, sorry. First, let me revise in terms of pi Given the complex number a + bi If a and b are BOTH POSITIVE\[\large A = \tan^{-1}\left|\frac{b}{a}\right|\] If a is negative and b is positive\[\large A =\pi -\tan^{-1}\left|\frac{b}{a}\right|\] If a and b are BOTH NEGATIVE\[\large A = \pi+\tan^{-1}\left|\frac{b}{a}\right|\] If a is positive and b is negative \[\large A = 2\pi -\tan^{-1}\left|\frac{b}{a}\right|\]

OpenStudy (anonymous):

would it be 2pi/3?

terenzreignz (terenzreignz):

One more thing. If you have the angle measure in degrees, you just multiply it by \[\large \frac{\pi}{180}\] To get it in radians. And you're correct, by the way, so what's your final answer?

OpenStudy (anonymous):

6 ( cos 2pi/3 + i sin 2pi/3)! Thank you so much!

terenzreignz (terenzreignz):

No problem :)

terenzreignz (terenzreignz):

Going off now ------ Terence out ;)